Write expression that gives the required unknown value.


1. If a = 8 and c = 12, find B.

Answer:


2. If a = 13 and B = 16°, find c.

Answer:


3. If A = 15° and c = 37, find a.

Answer:


1
Expert's answer
2022-04-13T18:09:41-0400


1_________________

"cosB = \frac{a}{c}\\
cosB = \frac{8}{12}\\
B = arccos(\frac{8}{12})"

2_________________

"cosB = \frac{a}{c} \\
c = \frac{a}{cosB}\\
c = \frac{13}{cos16°}"

3_________________

"sinA =\frac{a}{c}\\
a = c×sinA\\
a = 37×sin15°"